# A Stable Whitening Optimizer for Efficient Neural Network Training

### v2

@@ Abstract
In this work, we take an experimentally grounded look at neural network optimization. Building on the Shampoo family of algorithms, we identify and alleviate three key issues, resulting in the proposed SPlus method. First, we find that naive Shampoo is prone to divergence when matrix-inverses are cached for long periods. We introduce an alternate bounded update combining a historical eigenbasis with instantaneous normalization, resulting in across-the-board stability and significantly lower computational requirements. Second, we adapt a shape-aware scaling to enable learning rate transfer across network width. Third, we find that high learning rates result in large parameter noise, and propose a simple iterate-averaging scheme which unblocks faster learning. To properly confirm these findings, we introduce a pointed Transformer training benchmark, considering three objectives (language modelling, image classification, and diffusion modelling) across different stages of training. On average, SPlus is able to reach the validation performance of Adam within $44\%$ of the gradient steps and $62\%$ of the wallclock time.
@@ 1 Introduction
In experiments, SPlus achieves the fastest loss decrease among a wide suite of previous optimizers. SPlus matches Adam within $\mathbf{{\color[rgb]{0.82421875,0.078125,0.09375}\sim 44\%}}$ of the gradient steps, and within $\mathbf{{\color[rgb]{0.82421875,0.078125,0.09375}\sim 62\%}}$ of the wallclock time. SPlus does not introduce any critical hyperparameters to tune, and does not require additional forward/backward passes. We hope that this effectiveness and simplicity enables the community to easily adopt SPlus to their existing training settings.
@@ 3 Background and Preliminaries
Gradient descent methods can be seen as following the steepest descent direction under some metric. Naive gradient descent implicitly assumes a Euclidean distance metric over parameters, in which case the update vector is simply the scaled gradient itself:
$u=\operatorname*{arg\,min}_{\Delta\theta}\;\underbrace{\;g^{T}\Delta\theta\;}_{\text{Improvement}}+\underbrace{\alpha||\Delta\theta||^{2}}_{\text{Distance Penalty}}\quad=\quad\alpha g.$ (1)
However, it is often helpful to impose other metrics. For example, certain parameters may be more sensitive to second-order changes, and thus should be assigned a larger penalty. We can generally express distance using a Riemannian metric $M$, which is a symmetric positive-definite matrix of shape $\mathbb{R}^{dim(\theta)\times dim(\theta)}$. Under $M$, the distance of an update can be expressed as the matrix product:
$|\Delta\theta|^{2}_{M}=\Delta\theta^{T}M\Delta\theta.$ (2)
@@
$u=\operatorname*{arg\,min}_{\Delta\theta}\;\underbrace{\;g^{T}\Delta\theta\;}_{\text{Improvement}}+\underbrace{\Delta\theta^{T}M\Delta\theta}_{\text{Distance Penalty}}\quad=\quad M^{-1}g.$ (3)
which, in the case of Euclidean distance represented by the identity metric, reduces to Equation 1.
@@
What about the Hessian or the Fisher? While potentially effective, these metrics are often expensive to compute as they require additional information outside the standard gradient. We refer to detailed discussions in [34, 31, 48, 21], as well as a brief overview in Section A.4.
@@
In Shampoo [15] (Algorithm 1), the above techniques are utilized to derive an efficient update. The factor matrices of the whitening metric can be directly calculated from the matrix-shaped gradients:
$M^{2}=E[gg^{T}]=(L\otimes R)\quad\leftrightarrow\quad L=E[G^{T}G]\qquad R=E[GG^{T}]$ (8)
@@
$u=M^{-1}g\quad\leftrightarrow\quad U=L^{-1/2}GR^{-1/2}.$ (9)
In practice, matrix inversion is slow, so the above inversion is performed only every $N$ steps, and the results are cached until recomputation. This caching can result in unstable training (Figure 2), and is one of the issues we will discuss in the following sections.
@@ 4.1 To reliably prevent divergence, utilize instant-sign normalization
Plotted above are loss curves on language modelling, sweeping over learning rate between $(0.0001,0.000215,0.000464,0.001)$ and cache duration between $(5,10,25,100,500)$. SPlus is significantly more robust to hyperparameters than Shampoo. This robustness is crucial for improving practical training speed – while Shampoo diverges, SPlus remains stable when caching for $>100$ gradient steps, enabling a faster wall-clock performance than Adam.
As shown in Figure 2, we find that naive Shampoo is prone to divergence. When examining a range of learning rates and matrix-inversion frequencies, we find that Shampoo diverges in $>50\%$ of the trials. Notably, Shampoo always diverges when matrix-inversion is cached for over 25 gradient steps.
@@
$U_{Shampoo}=(Q_{L}\Lambda_{L}^{-1/2}Q_{L}^{T})\;G\;(Q_{R}\Lambda_{R}^{-1/2}Q_{R}^{T})$ (10)
@@
Instant-sign normalization has a hard bound preventing divergence. As $Q_{L}/Q_{R}$ are orthonormal, and the inner sign-matrix contains only $1$ or $-1$, the resulting update will satisfy:
$||U||_{spectral}\leq||U||_{frob}=\sqrt{nm}\qquad\text{and}\qquad||U||_{\infty}\leq\max(\sqrt{m},\sqrt{n}).$ (12)
@@
$U=\Lambda_{L}^{-1/2}G\Lambda_{R}^{-1/2}\quad\leftrightarrow\quad u=(\lambda_{L}\otimes\lambda_{R})^{-1/2}\;g$ (13)
where notably, the diagonal component of $(\lambda_{L}\otimes\lambda_{R})$ is not fully expressive due to being constructed out of a Kronecker product. This notion is studied in SOAP [52], who note that Shampoo is equivalent to a rank-1 Adam approximation in a rotated eigenbasis. Their proposed method alleviates this issue with an additional elementwise normalization matrix for each parameter. In contrast, instant-sign normalization does not require additional parameters in memory.
As shown in Figure 2 (left), the instant-sign normalization of SPlus eliminates divergence across the board. Crucially, SPlus allows for the matrix-inversion to be cached for significantly longer intervals without collapse. Prior works on Shampoo perform recomputation every 10 steps [46, 15], and our empirical findings support that this frequency is needed to prevent divergence. In contrast, SPlus remains stable even when results are cached for over 100 steps. Utilizing this stability, SPlus can be run at a speed which outperforms Adam in reaching an equivalent validation loss not only in gradient steps, but also in wall-clock time (Figure 2, right).
@@ 4.2 To standardize learning rate across network widths, use symmetric shape-aware scaling
Learning rate is often the first hyperparameter to tune due to its outsized impact on performance. Recent works have shown that for SGD and Adam, it is possible to naturally parameterize updates such that the optimal learning rate remains constant even as network width is adjusted [53, 54]. As shown in Figure 3, neither Shampoo nor the instant-sign update above display the correct learning-rate transfer across widths. In this section, we derive a simple adjustment to introduce learning-rate transfer to our setting as well.
@@
for any intermediate activation vector $\text{for any intermediate activation vector $x$:}\quad\sqrt{\dfrac{1}{k}\sum_{i=0}^{k}(\Delta x_{i})^{2}}=\mathbb{O}(1).$ (14)
To achieve the above property, the norm of the update to each dense layer of shape $U=\mathbb{R}^{m\times n}$ must be properly scaled. In SPlus, Equation 12 states that the Frobenius norm of a raw instant-sign update is $\sqrt{nm}$. Thus, when considering a network with $c$-times larger width, one should divide the update by a factor of $c$. We implement this scaling without a reference width by introducing a per-layer scaling factor of $2/(m+n)$:
[…]
As shown in Figure 3, the shape-aware scaling factor enables a natural parametrization where learning rates transfer between network width. This property is especially useful for tuning, and enables a more robust default learning rate for the optimizer.
Our scaling factor is different than the "spectral" scaling of $1/m$ introduced in [53], and this is by design. We found that spectral scaling was harmful in the MLP block. Consider a $(256,1024)$ layer and a $(1024,256)$ layer. Under spectral scaling, the first layer would have a $4x$ larger per-parameter learning rate than the second. In contrast, our method is symmetric and assigns the same learning rate to both layers (as Adam does) while properly normalizing when both widths are increased. We made this design choice purely on empirical findings – our average scaling outperformed the spectral scaling in all cases, see Figure 7 – and provide a brief discussion in Section A.3.
@@ 4.3 To reduce parameter noise, make use of iterate averaging
Figure 4 highlights the benefits of simple iterate averaging. Across the board, evaluating at the exponentially averaged parameters achieves a lower validation loss. Note that series of live parameters in the averaged and non-averaged cases are equivalent. The averaged parameters more closely reveal the “true” learning progress of utilizing a higher learning rate, which is otherwise obscured by parameter noise causing an increase in validation loss.
@@
        $\tilde{L}^{-1/2}\leftarrow invsqrt(L)$
        $\tilde{R}^{-1/2}\leftarrow invsqrt(R)$
@@
     $U\leftarrow\tilde{L}^{-1/2}\bar{G}\tilde{R}^{-1/2}$
@@ 5 How does SPlus compare to prior optimizers?
The specific Transformer architecture is adapted from GPT-2 [8]. Layer normalization terms are applied pre-attention and pre-MLP. We remove bias terms from the network. Each objective also includes different input/output heads – a token embedding and logit predictor for language modelling, a patch embedder and patch output for diffusion modelling, and a patch embedding and class predictor for image classification. We use a momentum of $0.9$ when applicable, a linear warmup of 200 steps, and a weight decay of $0.1$. We train in bfloat16. See the provided code for further details.
@@
Figure 6: SPlus outperforms prior methods in both gradient steps and wallclock time, and matches Adam performance within $\mathbf{44\%}$ of the gradient steps. Learning rates are swept independently for each method. We examine the training of a 160M-parameter Transformer with a batch size of 1024, and a sequence/patch length of 256. Results are averaged starting from three base checkpoints, and the full results are in Table 1 of the Appendix.
To thoroughly compare between optimizer types, we consider the performance across different stages of training. Concretely, we construct base checkpoints by with Adam, and saving checkpoints at fixed intervals (initialization, ten thousand, and fifty thousand steps). We then evaluate each optimizer on the three checkpoints, training for ten thousand additional gradient steps. Learning rate is swept independently for each optimizer type, along a resolution of $10^{1/3}$, e.g. $(0.0001,0.000215,0.000464,0.001,...)$. Final performance is reported as validation loss after this procedure, measured on a fixed held-out validation set. The same random seed and data order are used in each run.
As loss scales vary per objective, we focus on steps-to-Adam as the main metric. We record the fraction of gradient steps and fraction of wallclock time required to match the performance of Adam on the task, measured via validation loss. Wall-clock results are machine-specific; we run all experiments on the same set of 32 TPUv3 pods, a typical run takes half a day. Results are reported utilizing the best-performing learning rate for each optimizer.
@@
  • Schedule-Free Adam [9], which replaces traditional momentum with a set of live parameters, and a set of slow parameters. Gradients are evaluated at a linear interpolation of the two.
@@
  • Shampoo [15], as described in the background section. We do not use learning-rate grafting. Matrix inversion is performed every 10 steps (and otherwise diverges).
@@
  • Muon [18], which performs an orthogonalization procedure on each gradient via Newton-Schulz iteration. The orthogonalization is performed at every update, without historical information.
As displayed in Figure 6, SPlus is able to outperform prior methods across the board in both gradient steps and wallclock time. We find that a well-tuned Adam is a hard baseline to beat. For example, we were unable to match the performance of Adam with Sophia (a similar finding was reported in [57]) or with Muon for LLM training. Shampoo was especially unstable (as discussed in earlier sections), and we find that when divergence does not occur, Shampoo training curves roughly match those of SOAP, as also reported in [52]. In terms of wall-clock performance, a strong contender is Schedule-Free Adam, which does not perform any matrix-based computation and only utilizes elementwise operators. We did not heavily tune the matrix-inversion frequency for SPlus or SOAP, and assign it a default value of $100$ – additional tuning could further improve wall-clock efficiency.
@@ 5.1 What implementation details matter for SPlus?
We only apply the SPlus update on two-dimensional dense layers, which in the case of a Transformer, composes a majority of the backbone (the exception being LayerNorm scale parameters). We also do not apply SPlus to the domain-specific input and output layers – e.g. the token embedding, the classification head, and the convolutional patch layers. For these nonstandard parameters, we simply set the update as the sign of the momentum values. Additionally, for nonstandard parameters where the shape-dependent scaling term of Equation 15 is undefined, we use a fixed constant scaling ($0.001$ in our experiments). We found that this constant is not sensitive to even $10$x or $0.1$x pertubations, and does not need to be tuned.
[…]

### v3

@@ Abstract
In this work, we take an experimentally grounded look at neural network optimization. Building on the Shampoo family of algorithms, we identify and alleviate three key issues, resulting in the proposed SPlus method. First, we find that naive Shampoo is prone to divergence when matrix-inverses are cached for long periods. We introduce an alternate bounded update combining a historical eigenbasis with instantaneous normalization, resulting in across-the-board stability and significantly lower computational requirements. Second, we adapt a shape-aware scaling to enable learning rate transfer across network width. Third, we find that high learning rates result in large parameter noise, and propose a simple iterate-averaging scheme which unblocks faster learning. To properly confirm these findings, we introduce a pointed Transformer training benchmark, considering three objectives (language modelling, image classification, and diffusion modelling) across different stages of training. On average, SPlus is able to reach the validation performance of Adam within $44-58\%$ of the gradient steps and $62-83\%$ of the wallclock time.
@@ 1 Introduction
In experiments, SPlus achieves the fastest loss decrease among a wide suite of previous optimizers. SPlus matches Adam within $\mathbf{{\color[rgb]{0.82421875,0.078125,0.09375}\sim 44-58\%}}$ of the gradient steps, and within $\mathbf{{\color[rgb]{0.82421875,0.078125,0.09375}\sim 62-83\%}}$ of the wallclock time. SPlus does not introduce any critical hyperparameters to tune, and does not require additional forward/backward passes. We hope that this effectiveness and simplicity enables the community to easily adopt SPlus to their existing training settings.
@@ 3 Background and Preliminaries
Gradient descent methods can be seen as following the steepest descent direction under some metric. Naive gradient descent implicitly assumes a Euclidean distance metric over parameters, in which case the update vector is simply the scaled gradient $g=\nabla_{\theta}L(\theta,x)$ itself:
[…]
However, it is often helpful to impose other metrics. For example, certain parameters may be more sensitive to second-order changes, and thus should be assigned a larger penalty. We can generally express distance using a Riemannian metric tensor $M$, which is a symmetric positive-definite matrix of shape $\mathbb{R}^{dim(\theta)\times dim(\theta)}$. Under $M$, the distance of an update can be expressed as the matrix product:
$\Delta\theta||^{2}_{M}=\Delta\theta^{T}M\Delta\theta.$ (2)
@@
$u=\operatorname*{arg\,min}_{\Delta\theta}\;\underbrace{\;g^{T}\Delta\theta\;}_{\text{Improvement}}+\underbrace{(1/2)\Delta\theta^{T}M\Delta\theta}_{\text{Distance Penalty}}\quad=\quad M^{-1}g.$ (3)
which, in the case of Euclidean distance represented by the identity metric, reduces to Equation˜1.
@@
What about the Hessian or the Fisher? While potentially effective, these metrics are often expensive to compute as they require additional information outside the standard gradient. We refer to detailed discussions in [34, 31, 49, 21], as well as a brief overview in Section˜A.4.
@@
In Shampoo [15] (Algorithm˜1), the above techniques are utilized to derive an efficient update. The factor matrices of the whitening metric can be directly calculated from the matrix-shaped gradients:
$M=E[gg^{T}]^{-1/2}\quad\approx\quad(L\otimes R)\quad\text{where}\quad L=E[GG^{T}]^{-1/4}\qquad R=E[G^{T}G]^{-1/4}$ (8)
@@
$u=M^{-1}g\quad\approx\quad U=L^{-1/4}GR^{-1/4}.$ (9)
In practice, matrix inversion is slow, so the above inversion is performed only every $N$ steps, and the results are cached until recomputation. This caching can result in unstable training (Figure˜2), and is one of the issues we will discuss in the following sections.
@@ 4.1 To reliably prevent divergence, utilize instant-sign normalization
Plotted above are loss curves on language modelling, sweeping over learning rate between $(0.0001,0.000215,0.000464,0.001)$ and cache duration between $(5,10,25,100,500)$. SPlus is significantly more robust to hyperparameters than Shampoo. This robustness is crucial for improving practical training speed – in our setting. Shampoo diverges when caching for >100>100 gradient steps while SPlus remains stable, enabling a faster wall-clock performance than Adam.
As shown in Figure˜2, we find that naive Shampoo is prone to divergence. When examining a range of learning rates and matrix-inversion frequencies, we find that Shampoo diverges in $50\%$ of the trials in our setting. Notably, Shampoo regularly diverges when matrix-inversion is cached for over 25 gradient steps.
@@
$U_{Shampoo}=(Q_{L}\Lambda_{L}^{-1/4}Q_{L}^{T})\;G\;(Q_{R}\Lambda_{R}^{-1/4}Q_{R}^{T})$ (10)
@@
Instant-sign normalization has a hard bound preventing divergence. As $Q_{L}/Q_{R}$ are orthonormal, and the inner sign-matrix contains only $1$ or 1-1, the resulting update will satisfy:
[…]
@@
$U=\Lambda_{L}^{-1/4}G\Lambda_{R}^{-1/4}\quad\leftrightarrow\quad u=(\lambda_{L}\otimes\lambda_{R})^{-1/4}\;g$ (13)
where notably, the diagonal component of $(\lambda_{L}\otimes\lambda_{R})$ is not fully expressive due to being constructed out of a Kronecker product. This notion is studied in SOAP [53], who note that Shampoo (with a $1/2$ power) is equivalent to a rank-1 Adam approximation in a rotated eigenbasis. Their proposed method alleviates this issue with an additional elementwise normalization matrix for each parameter. In contrast, instant-sign normalization does not require additional parameters in memory.
As shown in Figure˜2 (left), the instant-sign normalization of SPlus eliminates divergence across the board. Crucially, SPlus allows for the matrix-inversion to be cached for significantly longer intervals without collapse. Prior works on Shampoo perform recomputation every 10 steps [47, 15], and our empirical findings support that this frequency is needed to prevent divergence. In contrast, SPlus remains stable even when results are cached for over 100 steps. Utilizing this stability, SPlus can be run at a speed which outperforms Adam in reaching an equivalent validation loss not only in gradient steps, but also in wall-clock time (Figure˜2, right).
@@ 4.2 To standardize learning rate across network widths, use symmetric shape-aware scaling
Learning rate is often the first hyperparameter to tune due to its outsized impact on performance. Recent works have shown that for SGD and Adam, it is possible to naturally parameterize updates such that the optimal learning rate remains constant even as network width is adjusted [54, 55]. As shown in Figure˜3, neither Shampoo nor the instant-sign update above display the correct learning-rate transfer across widths. In this section, we derive a simple adjustment to introduce learning-rate transfer to our setting as well.
@@
$\text{for any intermediate activation vector $x$:}\quad\sqrt{\dfrac{1}{k}\sum_{i=0}^{k}(\Delta x_{i})^{2}}=\mathbb{O}(1).$ (14)
To achieve the above property, the norm of the update to each dense layer of shape $U=\mathbb{R}^{m\times n}$ must be properly scaled. In SPlus, Equation˜12 states that the Frobenius norm of a raw instant-sign update is $\sqrt{nm}$. Thus, when considering a network with $c$-times larger width, one should divide the update by a factor of $c$. We implement this scaling without a reference width by introducing a per-layer scaling factor of $2/(m+n)$:
$U=Q_{L}sign(Q_{L}^{T}GQ_{R})Q_{R}^{T}\;*\;2\;/\;(m+n)$ (15)
As shown in Figure˜3, the shape-aware scaling factor enables a natural parametrization where learning rates transfer between network width. This property is especially useful for tuning, and enables a more robust default learning rate for the optimizer.
Our scaling factor is different than the "spectral" scaling of $1/m$ introduced in [54], and this is by design. We found that spectral scaling was harmful in the MLP block. Consider a $(256,1024)$ layer and a $(1024,256)$ layer. Under spectral scaling, the first layer would have a 4x larger per-parameter learning rate than the second. In contrast, our method is symmetric and assigns the same learning rate to both layers (as Adam does) while properly normalizing when both widths are increased. We made this design choice purely on empirical findings – our average scaling outperformed the spectral scaling in all cases, see Figure˜7 – and provide a brief discussion in Section˜A.3.
@@ 4.3 To reduce parameter noise, make use of iterate averaging
Figure˜4 highlights the benefits of simple iterate averaging. Across the board, evaluating at the exponentially averaged parameters achieves a lower validation loss. Note that series of live parameters in the averaged and non-averaged cases are equivalent. The averaged parameters more closely reveal the “true” learning progress of utilizing a higher learning rate, which is otherwise obscured by parameter noise causing an increase in validation loss.
@@
   $\tilde{L}^{-1/4}\leftarrow matpow(L,-1/4)$
   $\tilde{R}^{-1/4}\leftarrow marpow(R,-1/4)$
@@
  $U\leftarrow\tilde{L}^{-1/4}\bar{G}\tilde{R}^{-1/4}$
@@ 5 How does SPlus compare to prior optimizers?
The specific Transformer architecture is adapted from GPT-2 [8]. Layer normalization terms are applied pre-attention and pre-MLP. We remove bias terms from the network. Each objective also includes different input/output heads – a token embedding and logit predictor for language modelling, a patch embedder and patch output for diffusion modelling, and a patch embedding and class predictor for image classification. We use a momentum of $0.9$ when applicable, a linear warmup of 200 steps followed by a constant schedule, and a weight decay of $0.1$. We train in bfloat16. See the provided code for further details.
@@
Figure 6: SPlus outperforms prior methods in both gradient steps and wallclock time, and matches Adam performance within $\mathbf{44-58\%}$ of the gradient steps. Learning rates are swept independently for each method. We examine the training of a 160M-parameter Transformer with a batch size of 1024, and a sequence/patch length of 256. Results are averaged starting from three base checkpoints, and the full results are in Table˜3 of the Appendix.
To thoroughly compare between optimizer types, we consider the performance across different stages of training. Concretely, we construct base checkpoints by with Adam, and saving checkpoints at fixed intervals (initialization, ten thousand, and fifty thousand steps). We then evaluate each optimizer on the three checkpoints, training for ten thousand additional gradient steps. Learning rate is swept independently for each optimizer type, along a resolution of $10^{1/3}$, e.g. (0.0001,0.000215,0.000464,0.001,)(0.0001,0.000215,0.000464,0.001,...). Final performance is reported as validation loss after this procedure, measured on a fixed held-out validation set. The same random seed and data order are used in each run.
As loss scales vary per objective, we focus on steps-to-Adam as the main metric. We record the fraction of gradient steps and fraction of wallclock time required to match the performance of Adam on the task, measured via validation loss. Wall-clock results are machine-specific and should be seen as a rough estimate; we run all experiments on the same set of 32 TPUv3 pods, a typical run takes half a day. Results are reported utilizing the best-performing learning rate for each optimizer.
@@
  • Schedule-Free Adam [9], which replaces traditional momentum with a set of live and slow parameters. Gradients are evaluated at a linear interpolation of the two.
@@
  • Shampoo [15], as described in the background section. We do not use learning-rate grafting. Matrix inversion is performed every 10 steps (and otherwise diverges). We additionally consider Shampoo with a 1/2 power, following [36, 47].
@@
  • Muon [18], which performs an orthogonalization procedure on each gradient via Newton-Schulz iteration at every update, without historical information. As Muon utilizes a higher inherent learning rate, we use a weight decay of $0.001$.
As displayed in Figure˜6, SPlus is able to outperform prior methods across the board in both gradient steps and wallclock time. We find that a well-tuned Adam is a hard baseline to beat. For example, we were unable to match the performance of Adam with Sophia (a similar finding was reported in [58]) or with Muon for LLM training. Shampoo was especially unstable (as discussed in earlier sections), and we find that when divergence does not occur, Shampoo training curves roughly match those of SOAP, as also reported in [53]. In terms of wall-clock performance, a strong contender is Schedule-Free Adam, which does not perform any matrix-based computation and only utilizes elementwise operators. For the main table, we did not heavily tune the matrix-inversion frequency for SPlus or SOAP, and assign it a default value of 100100Table˜1 shows the effect of ablating this frequency.
@@ 5.1 What implementation details matter for SPlus?
For frequency $\rightarrow$ 1 5 10 25 100 250 500
Steps-to-Adam, SPlus 0.78 0.71 0.75 0.76 0.77 0.77 0.81
Steps-to-Adam, Shampoo 0.86 0.80 0.80 0.79 Div. Div. Div.
Time per Update, SPlus 7.39 1.79 1.09 0.70 0.48 0.44 0.43
Time per Update, Shampoo 7.41 1.78 1.09 0.67 0.42 0.41 0.40
Time per Update, Adam 0.34
Time-to-Adam, SPlus 16.21 3.34 2.05 1.20 0.74 0.65 0.67
Time-to-Adam, Shampoo 18.11 3.79 2.18 1.21 Div. Div. Div
Table 1: Eigendecomposition frequency trades off between performance and wallclock time. Each trial is run for 1000 gradient steps on the LLM-Init objective. Note that specific wallclock times are machine-dependent (we train on a v3-32 TPU) and may vary between systems.
Method LLM-Init LLM-10K LLM-50K Memory Usage
Adam 1.0 1.0 1.0 $3mn$
S.Free Adam > 1.0 0.532 0.49 $3mn$
Adam + Iterate Averaging 0.677 0.52 0.81 $4mn$
Shampoo Div. Div. 0.699 $2mn+2(m^{2}+n^{2})$
SPlus (No Averaging) 0.729 0.662 0.685 $2mn+2(m^{2}+n^{2})$
SPlus 0.487 0.422 0.348 $3mn+2(m^{2}+n^{2})$
SOAP 0.712 0.66 0.677 $3mn+2(m^{2}+n^{2})$
SOAP + Iterate Averaging 0.513 0.416 0.317 $4mn+2(m^{2}+n^{2})$
Table 2: Ablating the usage of iterate averaging vs. the underlying gradient transformation.
We only apply the SPlus update on two-dimensional dense layers, which in the case of a Transformer, composes a majority of the backbone (the exception being LayerNorm scale parameters). We also do not apply SPlus to the domain-specific input and output layers – e.g. the token embedding, the classification head, and the convolutional patch layers. For these nonstandard parameters, we simply set the update as the sign of the momentum values. Additionally, for nonstandard parameters where the shape-dependent scaling term of Equation˜15 is undefined, we use a fixed constant scaling ($0.001$ in our experiments). We found that this constant is not sensitive to even 1010x or 0.10.1x pertubations, and does not need to be tuned.
We perform the above experiments over a pod of 32 TPUv3 machines, and parameters are distributed in a fully-sharded data parallel [59] setup. The SPlus update is distributed among devices. Specifically, the per-step computations are calculated independently on each device as usual. However, for the eigendecomposition which occurs every N steps, we instead broadcast the $L$ and $R$ matrices evenly among devices. In parallel, each device then performs the eigendecomposition for its assigned matrices, then re-broadcasts the results. In this way, the most expensive step of the SPlus update can be reduced by a factor of $(1/\text{num devices})$.
To ablate the specific changes proposed in SPlus, we show in Table˜2 an ablation on the interaction between iterate-averaging and instant-sign normalization, focusing specifically on the LLM setting. Iterate averaging is consistently helpful regardless of the underlying gradient transformation. This trend may be a result of our usage of a constant learning rate; as discussed in [9, 17], iterate averaging can be seen as a stand-in for learning rate decay. The downside is the memory requirement for an additional set of parameters. We note that SOAP + Iterate Averaging is competetive and can outperform SPlus. One way to contrast these methods is that the instant-sign normalization used in SPlus is a memory-efficient approximation of the internal Adam optimizer used in SOAP, which otherwise requires an additional set of parameters in memory.